ted is ten times as old as his daughter sarah. in four years he will be 6 times as old as she how old is sarah?

Answer by scott8148(6628)

"ted is ten times as old as his daughter sarah" ___ t = 10s

"in four years he will be 6 times as old as she" ___ t + 4 = 6(s + 4)

substituting ___ (10s) + 4 = 6s + 24 ___ 4s = 20